An investor invested a total of $2,500 in two mutual funds. One earned a 4% profit while the other earned a 2% profit. If the investor's total profit was $88, how much was invested in each mutual fund?

Answer by checkley79(3341)

.04x+.02(2500-x)=88
.04x+50-.02x=88
.02x=88-50
.02x=38
x=38/.02
x=$1900 amount invested @ 4%.
2500-1900=$600 invested @ 2%.
Proof:
.04*1900+.02*600=88
76+12=88
88=88